The product of two consecutive positive even integers is one hundred sixty-eight. Find the two integers.

Answer by unlockmath(1688)

Hello,
Let's show the numbers as "x" and "x+2" then we can solve this as:
x(x+2)=168 Subtract 168 from both sides to get:
x^2+2x-168=0
Now factor this:
(x-12)(x+14)=0 Now solve x to get:
x=12
x=-14
Obviously the only positive number would be:
12 and the other number is 14
Make sense?
RJ
There's a book I wrote on the basics of math:
www.math-unlock.com